A piston cylinder contains gas initially at 3500 kPa with a volume of 0.03 cubic meter. The gas is compressed during a process where pV raise to 1.25 = C to a pressure of 8500 kPa. The heat transfer from the gas is 2.5 kJ. Determine the nonflow work in kj, neglecting changes in kinetic and potential energies.
1
Expert's answer
2020-10-08T15:10:52-0400

Initial pressure(p1)=3500x103 Pa

Initial Volume (V1)=0.03 m3

Final pressure(p2)=8500 x103 Pa

Gas undergoes a polytropic process given by pV1.25=constant

Therefore, for process 1-2 we can write,

"p_1 \times {V_1}^{1.25} = p_2 \times {V_2}^{1.25}"

"{V_2}^{1.25} =\frac { p_1 \times {V_1}^{1.25} }{p_2}" =0.01475 m3

The work done during poly tropic process is given by

"W=\frac {p_1 \times V_1-p_2 \times V_2}{n-1}"

"W=\frac {(3500 \times (0.03)-(8500)\times (0.01475)}{1.25-1}" =-81.5 kJ (-ve sign indicates work done on the gas)

Work done on the gas is W=81.5 kJ